OpenStudy (anonymous):

write a quadratic equation in the variable x having the given number as solutions. Type in standard form, ax^2+bx+c=0. Solution; 2, only solution.

OpenStudy (anonymous):

(x-2)*(x-2)=0 expand it

OpenStudy (anonymous):

x^2-4*x+4

OpenStudy (anonymous):

x^2-4*x+4=0

Parth (parthkohli):

Well, one solution: \(\Large \color{Black}{\Rightarrow (x + 2)^2 = 0 }\) \(\Large \color{Black}{\Rightarrow x^2 + 4x + 4 = 0 }\) Two solutions: \(\Large \color{Black}{\Rightarrow (x + 2)(x + 1) = 0 }\) \(\Large \color{Black}{\Rightarrow x(x + 1) + 2(x + 1) = 0 }\) \(\Large \color{Black}{\Rightarrow x^2 + x + 2x + 2 = 0 }\) \(\Large \color{Black}{\Rightarrow x^2 + 3x + 2 = 0 }\)

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

it is (x-2)^2=0 if you pick plus 2 then solution will be -2 not 2

OpenStudy (anonymous):

@ParthKohli it is (x-2)^2=0 if you pick plus 2 then solution will be -2 not 2

Parth (parthkohli):

Oh I thought we were talking about number of solutions haha

OpenStudy (anonymous):

Thats better than the answer I arrived at. ha

OpenStudy (anonymous):

x^2-4*x+4=0